# VLAP: Efficient Video-Language Alignment via Frame Prompting and Distilling for Video Question Answering

### v1

@@ Abstract
In this work, we propose an efficient Video-Language Alignment via Frame-Prompting and Distilling (VLAP) network. Our VLAP model addresses both efficient frame sampling and effective cross-modal alignment in a unified way. In our VLAP network, we design a new learnable question-aware Frame-Prompter together with a new cross-modal distillation (QFormer-Distiller) module. Pre-trained large image-language models have shown promising results on problems such as visual question answering. However, how to efficiently and effectively sample image frames when adapting pre-trained large image-language model to video-language alignment is still the major challenge. Compared with prior work, our VLAP model demonstrates the capability of selecting key frames with critical contents, thus improving the video-language alignment accuracy while reducing the inference latency (+3.3% on NExT-QA Temporal with 3.0$\times$ speed up). Overall, our VLAP network outperforms (e.g. +4.6% on STAR Interaction and +2.2% on STAR average with 3.0$\times$ speed up, ours 2-frames out-perform SeViLA 4-frames on VLEP with 4.2$\times$ speed up) the state-of-the-art methods on the video question-answering benchmarks.
@@ 1 Introduction
“If a picture is worth thousands of words, what is a video worth?” [32] Video watching is growing into a new social norm. Statistic shows YouTube has approximately $122$ million daily active users, based all over the world. Visitors spend on average $19$ minutes per day on YouTube. An average of close to 1 million hours of video are streamed by YouTube users each and every minute. As video data continue to grow through internet, video information retrieval becomes more and more demanding. Video data has tremendous capacity to store a vast variety of useful information. Compared to images, video understanding is difficult in the one extra temporal dimension. How to efficiently sample relevant frames from a video with the computing resource constraint remains a long-standing problem in video understanding research.
@@
Figure 1: Our efficient Vision-Language Alignment via Frame-Promper and distilling (VLAP) model, at core, contains two new modules: a question-aware Frame-Prompter and a cross-modal QFormer-Distiller. It learns to extract the most question-related frames while keeping the inference latency low.
@@
Recent advances in pre-trained large-scale language models [6, 42] have greatly boosted the performance of the vision-language models. Especially for image-language models [7, 1, 23, 46], many state-of-the-art works leverage pre-trained LLMs. These models  [9, 57] achieve excellent performance across visual-language tasks such as image captioning [9], visual question answering [57] and more. Inherently, many video-language models [62, 29] are built from these pre-trained image-language models. These image-based video-language models treat a video as a series of multi-channel images sampled randomly or uniformly [23]. While this strategy works well for short videos, for long videos or videos with non-uniform information distribution, random or uniform frame sampling may miss critical information. When it comes to video-language alignment, frame sampling efficiency and effectiveness goes hand-in-hand. One needs to not only reduce the number of sampled frames but also select frames that are most related to the input question. Previous work such as SeViLA[62] trains a separate keyframe localizer, which is not friendly for the real-time inference and introduce more parameters to the whole model. Besides video representation, cross-modality alignment while leveraging LLMs is another challenge. The critical problem lies in how to transfer video information to the LLM’s input domain.
@@
Figure 2: Model Overview. Our VLAP model includes 4 sub-modules: the visual encoder, question-aware Frame-Prompter (FP), QFormer-Distiller (QFD), and a LLM. We encode the video frames through a frozen visual encoder. Then we train the Teacher-QFormer using all the frame features. After that, we train the Student-QFormer and Frame-Prompter end-to-end. Unlike the Teacher-QFormer, our Student-QFormer is trained with masked frames features from a question-aware Frame-Prompter. Finally, the input question text and QFormer transformed visual features go through a frozen large language model to generate the answer.
@@
To address these challenges, we propose a new network, VLAP. Compared to the state-of-the-art video-language models [62, 9, 23], VLAP consists of a new question-aware Frame-Prompter together and a new QFormer-Distiller. To align image and language domains, we apply the cross-modal query-visual Transformer (QFormer) proposed in previous BLIP models [9, 23, 24] for cross-modal fusion. To improve the efficiency, we add a new distillation on top of the QFormer. We train our new Frame-Prompter and new QFormer-Distiller end-to-end. Our new question-aware Frame-Prompter learns to select the most relevant frames according to the input question text, and the QFormer-Distiller further reinforces the critical frames selection. To the best of our knowledge, we are the first to propose a Frame-Prompter and a distiller on top of the cross-modal alignment for video-language learning with pre-trained LLMs.
@@
We validate our VLAP model on the Video Question Answering benchmark datasets. This includes the NExT-QA [52], STAR[50], How2QA [27], TVQA [21] and VLEP [22]. Our work outperforms previous strong SOTA methods across all the benchmark while reducing inference latency. Comparing with SOTA video-language model SeViLA [62], we reduce significant amount ($50\%$) of training parameters and inference latency (1.354.2×1.35-4.2$\times$1.35 - 4.2 × speed up), while improving the accuracy. We also performed ablation study on our question-aware Frame-Prompter and QFormer-Distiller. As shown in Table 4, both the question-aware Frame-Prompter (FP) and the QFormer-Distiller (QFD) play critical roles in making our method effective.
To summarize, our main contribution is a new question-aware Frame-Prompter and QFormer-Distiller component (trained from end-to-end) to efficiently and effectively select frames for video-language alignment learning.
@@ 2.1 Visual-Language Alignment
Visual-Language Pre-training Vision-Language cross-modal pre-training has greatly improved over the past couple of years. Various network architectures and pre-training objectives have been proposed for different downstream tasks, including the dual-encoder architecture with image-text contrastive learning [38], the fusion-encoder architecture with image-text matching [25], and unified transformer architecture with masked language modeling [45]. These methods, along with others, focus on the ability to find image-text affinity [58], correlation [4], and/or completion [61], and need to pre-train the model end-to-end. To address the incompatibility with pre-trained unimodal models such as LLMs [6], recent works [23] proposed to train a QFormer to bridge the domain gap between two frozen pre-trained models. Inspired by its flexibility, more downstream tasks and applications have been proposed, including instruction-based image generation [51] and image question-answering [57].
While most of the previous work focus on image-text alignment, very few have discussed the extension to videos until most recently when temporal modeling starts to be included for better reasoning capabilities. HiTeA [59] jointly trains pairs in long and short view to capture temporal relations between moments and event. Smaug [28] introduces sparse image patch masks to reduce pre-training costs and improve cross-modal alignment. EgoVLPv2 [36] proposes cross-attention between the backbone encoders to improve both the pre-training efficiency and the downstream task performance. Yang et al. [56] models trajectory attention as a two-step attention operation to further emphasize word-related embeddings during the pre-training.
Our method leverages the pre-trained model during the video-language alignment, which naturally provides more capability to reasoning with less training cost.
@@
Due to the high cost of large video dataset collection, many works leverage successful pre-trained image models and transfer the knowledge to video task. Previous works such as  [3, 5, 16, 26] utilize a pretrained ViT [11] and aggregate the temporal image feature sequence using transformer block to adapt for video understanding task. For Video-Language task, many works turn to large-scale Vision-Language models as the starting point, such as CLIP [38], BLIP [24]. Many works choose to adapt a pre-trained CLIP model for text-to-video retrieval task, by either augmenting the frame-level spatial features with temporal information [53, 10], or manipulate the cross-modal similarity calculation to get better video-language alignment [31, 30, 12, 49]. Other works  [35, 37] focus on parameter efficient fine-tuning for video task by inserting trainable temporal modules into the pre-trained transformer architecture while keeping the rest of model frozen. Recent work SeViLA [62] proposes a language-aware frame localizer to sample relevant key frames from videos. It adopts a pre-trained Vision-Language model BLIP2 [23] as the frozen backbone and only the adapter QFormer is trained. In this paper, we propose a trainable question-aware Frame-Prompter and a QFormer-Distiller module. These techniques help to efficiently bridge the gap between image-language and video-language learning.
@@ 3 Method
Our VLAP model tackles the following challenges in large-scale Video-Language learning: how to sample question related frames and how to transfer visual information for pre-trained frozen LLMs. In this section, we describe our question-aware Frame-Prompter and QFormer-Distiller.
@@ 3.1 Model Architecture
As illustrated in Fig. 2, VLAP has four components: a pre-trained frozen large scale visual encoder $E_{v}$, a question-aware Frame-Prompter $F_{p}$, a QFormer-Distiller $Q_{d}$ (Querying Alignment Transformer [24, 23, 9] with distillation) to fuse and extract text-conditioned visual information, and a pre-trained frozen LLM. We train our Frame-Prompter, QFormer-Distiller and two other frozen modules end-to-end. Our training objective include a distillation loss $L_{\text{distill}}$ and a task loss $L_{\text{task}}$. For the multi-choice Video Question Answer task, the task loss is the cross-entropy loss.
3.2 Question-aware Frame-Prompter
Figure 3: Frame-Prompter. Here we show the details of our learnable Frame-Prompter. We design a learnable Frame-Prompter to sample the most text query-related frames. We encode the mean-pooled segment features. We then apply the Gumbel Softmax to compute the segment mask to guarantee the differentiability. The selected frames embedding then goes through the QFormer-Distiller. Here $B$ means batch size, $T$ means frame number, $N\times C$ means the frame feature sequences.
For the four-dimensional video data, it’s impractical and non-efficient to input all frames into a visual encoder model. Here we design a question-aware Frame-Prompter for efficient and effective frame sampling. It’s designed learn to attend to fewer but more question related frames. We start from the uniformly sampled video frames $\{f_{1},\dots,f_{T}\}$, $T$ is frame number. As shown in Fig. 3, these raw frames first go through the pre-trained visual encoder $E_{v}$,
$X=\{x_{i}|x_{i}=E_{v}(f_{i}),i\in[1,T]\},$ (1)
where $x_{i}$ is the visual feature extracted from raw frames and $T$ is the number of frames. We perform mean-pooling on the channels per-patch then encode the mean-pooled visual features onto an embedding using a fully-connected (FC) layer with a layer normalization (LN),
$\hat{x_{i}}=W_{2}*\operatorname{ReLU}(\mathrm{LN}(W_{1}*\operatorname{Mean}(x_{i}))),$ (2)
@@ 3.2 Question-aware Frame-Prompter
Our Frame-Prompter is a differentiable frame selecter. The differentiability is achieved through the Gumbel-Softmax [18]. We first convert a segment of concatenated frames feature $s_{k}=[x_{i},\dots,x_{S}]$ into a categorical distribution $\pi$ through the softmax operation,
$\pi=\left\{p_{k}\mid p_{k}=\frac{\exp(s_{k})}{\sum_{j=1}^{T/S}\exp(x_{j})},k,j\in[1,T/S]\right\},$ (3)
@@
[…]
@@
$\frac{\exp((\log p_{k}+g_{k})/\tau)}{\sum_{j=1}^{T/S}\exp((\log p_{j}+g_{j})/\tau)},k,j\in[1,T/S],$ (5)
where $\tau$ is a tunable temperature hyper-parameter.
@@ 3.3 Cross-Modal Distillation
In this section, we present our QFormer-Distiller, designed to further enforce the text conditioned visual information extraction. Video-language alignment performance, unlike image-language alignment, is closely related the selected frames. Meanwhile, to leverage powerful pre-trained LLMs, we need to transform the selected visual information to the LLM’s input domain. We adopt the QFormer proposed in the BLIP models [24, 23, 9] for the cross-modal transformation learning.
The original QFormer is designed for image-text fusion, and we add a distillation module to make it efficient for video-text fusion. Like the traditional distillation, our QFormer-Distiller includes a teacher-QFormer and a student-QFormer. We train the teacher-QFormer first (without the student-QFormer) on a wider receptive field. Then we train our student-QFormer to narrow down the most important frames. The student-teacher learning mechanism further encourages both the QFormer and the Frame-Prompter to learn to attend to the most relevant visual information given the input question text.
@@
$X^{\prime}=\operatorname{CrossAttn}(E_{t}(X_{q}),\operatorname{SelfAttn}(E_{v}(X_{v}),\mathbf{q})),$ (6)
where $E_{t}$ is a text tokenizer, $E_{v}$ is the visual tokenizer, $\mathbf{q}$ is the learnable query token. We utilize a decoder $D$ to transform the student-QFormer output, ensuring consistency and recoverability to the teacher’s feature. Then the distillation objective is defined by:
$L_{\text{distill}}=\operatorname{MSE}(D(X^{\prime}_{\text{student}}),X^{\prime}_{\text{teacher}}).$ (7)
We carefully design the decoder to be a simple Fully Connected (FC) layer with a layer normalization. We show through an ablation study (Sec. 5.3.2) that this combination works the best.
4 Experiment Setup
Here we introduce our implementation details, benchmark dataset and the baseline methods.
4.1 Implementation Details

### v3

# ViLA: Efficient Video-Language Alignment for Video Question Answering

@@ Abstract
In this work, we propose an efficient Video-Language Alignment (ViLA) network. Our ViLA model addresses both efficient frame sampling and effective cross-modal alignment in a unified way. In our ViLA network, we design a new learnable text-guided Frame-Prompter together with a cross-modal distillation (QFormer-Distiller) module. Pre-trained large image-language models have shown promising results on problems such as visual question answering (VQA). However, how to efficiently and effectively sample video frames when adapting pre-trained large image-language model to video-language alignment is still the major challenge. Compared with prior work, our ViLA model demonstrates the capability of selecting key frames with critical contents, thus improving the video-language alignment accuracy while reducing the inference latency (+3.3% on NExT-QA Temporal with 3.0$\times$ speed up). Overall, our ViLA network outperforms the state-of-the-art methods on the video question-answering benchmarks: +4.6% on STAR Interaction, +2.2% on STAR average with 3.0$\times$ speed up, ours 2-frames out-perform SeViLA 4-frames on the VLEP dataset with 4.2$\times$ speed-up.
@@ 1 Introduction
“If a picture is worth thousands of words, what is a video worth?” [36] Video watching has become a new social norm. Statistics show YouTube has approximately $122$ million daily active users, based all over the world. Visitors spend on average $19$ minutes per day on YouTube. An average of close to 1 million hours of video are streamed by YouTube users each and every minute. As video data continue to grow through internet viewing, video information retrieval becomes more demanding. Video data has tremendous capacity to store a vast variety of useful information. Compared to image question answering (Q&A) problem, video QA is more challenging due to one extra temporal dimension. How to efficiently sample relevant frames from a video with the computing resource constraint remains a long-standing problem in video QA research.
@@
Figure 1: Our efficient Vision-Language Alignment (ViLA) model via Frame-Promper and distilling contains two new modules: a text-guided Frame-Prompter and a cross-modal QFormer-Distiller. It learns to extract the most question-related frames while keeping the inference latency low.
@@
Recent advances in pre-trained large-scale language models [6, 46] have greatly boosted the performance of the vision-language models, especially on cross-modality tasks. Many state-of-the-art image-language models [8, 1, 27, 49] leverage pre-trained LLMs. These models  [10, 60] achieve excellent performance on visual-language tasks such as image captioning [10], visual question answering [60] and more. Inherently, many video-language models [65, 33] are built from these pre-trained image-language models. These image-based video-language models treat a video as a series of multi-channel images sampled randomly or uniformly [27]. While this strategy works well for short videos, for long videos or videos with non-uniform information distribution, random or uniform frame sampling may miss critical information. When it comes to video-language alignment, frame sampling efficiency and effectiveness go hand-in-hand. One needs to not only reduce the number of sampled frames but also select frames that are most related to the input question. Previous work such as SeViLA[65] trains a separate keyframe localizer, which is not friendly for the real-time inference and introduces more parameters to the whole model. Besides video representation, cross-modality alignment while leveraging LLMs is another challenge. The critical problem lies in how to efficiently transfer video information to the LLM’s input domain.
@@
Figure 2: Model Overview. Our ViLA model includes 4 sub-modules: the visual encoder, text-supervised Frame-Prompter (FP), QFormer-Distiller (QFD), and a LLM. We encode the video frames through a frozen visual encoder. Then we train the Teacher-QFormer using all the frame features. After that, we train the Student-QFormer and Frame-Prompter end-to-end. Unlike the Teacher-QFormer, our Student-QFormer is trained with masked frames features from a text-supervised Frame-Prompter. Finally, the input question text and QFormer transformed visual features go through a frozen large language model to generate the answer. Our network supports both leveraging LLM through proper visual prompting without affecting the original LLM (Frozen) ability on language tasks and fineuning LLMs(LoRA) simultaneously to get optimal performance on specific tasks.
@@
Main Contributions: To address these challenges, we propose a new network, ViLA. Compared to the state-of-the-art video-language models [65, 10, 27], ViLA consists of a new Frame-Prompter together with a QFormer-Distiller. Our Frame-Prompter learns to select the most important frames influenced by the corresponding question text and and supervised by the VQA loss. Meanwhile, the Frame-Prompter is meticulously designed to keep lightweight so as to be efficient. To effectively and efficiently transfer video information to LLM input domain, we add a new distillation on top of the QFormer, named QFormer-Distiller. The QFormer is the cross-modal query-visual Transformer proposed in previous BLIP models [10, 27, 28] for cross-modal fusion. We train our Frame-Prompter and QFormer-Distiller end-to-end. The cross-modal temporal distiller teaches a smaller (i.e. fewer frames) QFormer. To the best of our knowledge, this work is the first to propose a Frame-Prompter and a distiller on top of the cross-modal alignment for video-language learning with pre-trained LLMs.
@@
We validate our ViLA model on the Video Question Answering benchmark datasets. This includes the NExT-QA [56], STAR[54], How2QA [31], TVQA [25] and VLEP [26]. Our work outperforms previous strong SOTA methods across all the benchmarks, while reducing inference latency. Comparing with SOTA video-language model SeViLA [65], we reduce significant amount ($50\%$) of training parameters and inference latency (1.354.2×1.35-4.2$\times$1.35 - 4.2 × speed up), while improving the accuracy. We also performed ablation study on our text-guided Frame-Prompter and QFormer-Distiller. As shown in Table 4, both the text-guided Frame-Prompter and the QFormer-Distiller play critical roles in making our method effective.
To sum up, the key novelty is a new text-guided Frame-Prompter and question-relevant QFormer-Distiller (trained from end-to-end). The Frame-Prompter enhances the efficiency, while the later bolsters the effectiveness. Together they optimize the selection of frames for video-language alignment learning.
@@ 2.1 Visual-Language Alignment
Visual-Language Pre-training
Vision-Language cross-modal pre-training has greatly improved over the past couple of years. Various network architectures and pre-training objectives have been proposed for different downstream tasks, including the dual-encoder architecture with image-text contrastive learning [42], the fusion-encoder architecture with image-text matching [29], and unified transformer architecture with masked language modeling [50]. These methods, along with others, focus on the ability to find image-text affinity [61], correlation [4], and/or completion [64], and need to pre-train the model end-to-end. To address the incompatibility with pre-trained unimodal models such as LLMs [6], recent works [27] proposed to train a QFormer to bridge the domain gap between two frozen pre-trained models. Inspired by its flexibility, more downstream tasks and applications have been proposed, including instruction-based image generation [55] and image question-answering [60].
While most of the previous work focus on image-text alignment, very few have discussed the extension to videos until most recently when temporal modeling starts to be included for better reasoning capabilities. HiTeA [62] jointly trains pairs in long and short view to capture temporal relations between moments and event. Smaug [32] introduces sparse image patch masks to reduce pre-training costs and improve cross-modal alignment. EgoVLPv2 [40] proposes cross-attention between the backbone encoders to improve both the pre-training efficiency and the downstream task performance.
Our method can leverages the pre-trained model during the video-language alignment, which naturally provides reasoning capability with less training cost.
@@
Due to the high cost of large video dataset collection, many works leverage successful pre-trained image models and transfer the knowledge to video task. Previous works such as  [3, 5, 18, 30] utilize a pretrained ViT [12] and aggregate the temporal image feature sequence using transformer block to adapt for video understanding task. For Video-Language task, many works turn to large-scale Vision-Language models as the starting point, such as CLIP [42], BLIP [28]. Many works choose to adapt a pre-trained CLIP model for text-to-video retrieval task, by either augmenting the frame-level spatial features with temporal information [57, 11], or manipulate the cross-modal similarity calculation to get better video-language alignment [35, 34, 14, 53]. Other works  [39, 41] focus on parameter efficient fine-tuning for video task by inserting trainable temporal modules into the pre-trained transformer architecture while keeping the rest of model frozen. Recent work SeViLA [65] proposes a language-aware frame localizer to sample relevant key frames from videos. In this paper, we propose a trainable text-guided Frame-Prompter and a QFormer-Distiller module, which help focus more on the important temporal and spatial information from the 2D frames. These techniques help to efficiently bridge the gap between image-language and video-language learning.
@@ 2.2 Knowledge Distillation
2.3 Frame Selection for Video QA
Early approaches in Video QA relied heavily on dense sampling methods, where frames are extracted at a fixed interval throughout the video. While straightforward, such methods can lead to excessive computational costs and memory requirements without significantly improving performance. Zhang et al. [68] proposed a more selective strategy, using attention mechanisms to identify key frames that are more likely to contain information relevant to the question. Following this, adaptive frame sampling methods  [13, 7] have gained popularity. These methods aim to dynamically select frames based on the content’s relevance to the question, thus optimizing the trade-off between computational efficiency and answer accuracy [13].
@@ 3 Method
Our ViLA model tackles the following challenges in large-scale Video-Language learning: how to sample question related frames and how to efficiently transfer video information for pre-trained frozen LLMs.
@@ 3.1 Model Architecture
As illustrated in Fig. 2, ViLA has four components: a pre-trained frozen large scale visual encoder $E_{v}$, a Frame-Prompter $F_{p}$, a QFormer-Distiller $Q_{d}$ (Querying Alignment Transformer [28, 27, 10] with distillation) to fuse and extract text-conditioned visual information, and a pre-trained frozen LLM. We train our Frame-Prompter, QFormer-Distiller and two other frozen modules end-to-end. Our training objective includes a distillation loss $L_{\text{distill}}$ and a task loss $L_{\text{vqa}}$. For the multi-choice Video Question Answer task, the task loss is the cross-entropy.
3.2 Text-guided Frame-Prompter Learning
Figure 3: Text-guided Frame-Prompter. Here we show the details of our learnable text-guided Frame-Prompter. We design a learnable Frame-Prompter to sample the most text query-related frames, with two design choises (a and b). We choose design (a) for diversified temporal sampling. We first encode the mean-pooled segment features. We then apply the Gumbel Softmax to compute the segment mask to guarantee the differentiability. The selected frames embedding then goes through the QFormer-Distiller. Here $B$ means batch size, $T$ means frame number, $N\times C$ means the frame feature sequences. The Frame-Prompter is learned with the text-supervised gradient. When VQA loss is applied, the input question text-related gradient further flows to the Frame-Prompter. The question text-related gradient guides the Frame-Prompter to select the most critical frames.
For the four-dimensional video data, it’s impractical and non-efficient to input all frames into a visual encoder model. Here we design a text-guided Frame-Prompter for efficient and effective frame sampling. It is designed to learn attending to fewer but more question-related frames. We start from the uniformly sampled video frames $\{f_{1},\dots,f_{T}\}$, $T$ is frame number. As shown in Fig. 3, these raw frames first go through the pre-trained visual encoder $E_{v}$,
$X=\{x_{i}|x_{i}=E_{v}(f_{i}),i\in[1,T]\},\vspace{-2mm}$ (1)
where $x_{i}$ is the visual feature extracted from raw frames and $T$ is the number of frames. We perform mean-pooling on the channels per-patch. The Frame-Prompter input shape is $(B,T,N,C)$, B batch size, T temporal, $(N,C)$ as the frame feature dimension. Mean-pool averages over the $C$ channel. After mean-pool, the feature dimension becomes $(B,T,N)$. Then we divide the feature into S segments $(B,S,T//S\cdot N)$, with $(T//S\cdot N)$ as the segment feature dimension. Then the segmented features goes through FC layer to project to $(B,S,T//S)$ dimension, ready for the gumbel-softmax computation. For free-form frame sampling Fig. 3(b), after mean-pool, we use FC layer to transfer $(B,T,N)$ to $(B,T,2)$. Then we encode the mean-pooled visual features onto an embedding using a fully-connected (FC) layer with a layer normalization (LN),
$\hat{x_{i}}=W_{2}*\operatorname{ReLU}(\mathrm{LN}(W_{1}*\operatorname{Mean}(x_{i}))),\vspace{-2mm}$ (2)
@@ 3.2 Text-guided Frame-Prompter Learning
Our Frame-Prompter is a differentiable frame selecter. This allows the text-supervised gradients from the QFormer to guide the learning our frame selecter via backpropagation. The differentiability is achieved through the Gumbel-Softmax [21]. We have two choices (shown in Fig. 3 a and b) for frame selection before the Frame-Prompter: a) strategic sampling (like: segmented uniform sampling) and b) free-form sampling. We choose option (a) to force diverse temporal sampling. We first convert a segment of concatenated frames feature $s_{k}=[\hat{x_{i}},\dots,\hat{x}_{T//S}]$ into a categorical distribution $\pi$ through the softmax operation,
$\pi=\left\{p_{k}\mid p_{k}=\frac{\exp(s_{k})}{\sum_{j=1}^{T/S}\exp(s_{j})},k,j\in[1,T/S]\right\},\vspace{-2mm}$ (3)
@@
$M_{S}=\mathcal{H}\left(\arg\max_{k}\left[g_{k}+\log p_{k}\right]\right)k\in[1,T/S],\vspace{-2mm}$ (4)
@@
$\frac{\exp((\log p_{k}+g_{k})/\tau)}{\sum_{j=1}^{T/S}\exp((\log p_{j}+g_{j})/\tau)},k,j\in[1,T/S],\vspace{-2mm}$ (5)
where $\tau$ is a tunable temperature. Full mask is $M=\{M_{i},\cdots,M_{T//S}\}$. We apply mask $M$ to input frames and then conduct CrossAttention to obtain text guidance by
$X_{LLM}=\operatorname{CrossAttn}(X\cdot M,X_{t}),$ (6)
where $X_{t}$ is the text information. This CrossAttention can help the Frame-Prompter learn to choose rich text frames. Finally, we use the task loss below to supervise the learning and choose the frame selection to answer our question:
$L_{\text{vqa}}=\operatorname{MSE}(LLM(X_{LLM},X_{t}),QA_{\text{answer}}).\vspace{-2mm}$ (7)
@@ 3.3 Cross-Modal Distillation
In this section, we present in detail our cross-modal distillation module, designed to selectively transfer video information ready for LLM consumption. Video-language alignment performance, unlike image-language alignment, is closely related the selected frames. Meanwhile, to leverage powerful pre-trained LLMs, we need to transform the selected visual information to the LLM’s input domain. We adopt the QFormer proposed in the BLIP models [28, 27, 10] for the cross-modal transformation learning.
The original QFormer is designed for image-text fusion, and we add a distillation module to make it efficient for video-text fusion. Like the traditional distillation, our QFormer-Distiller includes a teacher-QFormer and a student-QFormer. We train the teacher-QFormer first (without the student-QFormer) on a wider receptive field. The student-teacher learning mechanism further encourages both the QFormer and the Frame-Prompter to learn to attend to the most relevant visual information given the input question text. We study how QFormer-Distiller affects Frame-Prompter (Sec.  4.3.1). And we find our FP+QFD 4-frames(Acc.74.4%) model’s key-frame selection overlaps 92.3% with the key-frame selected by the FP 8-frames(Acc.74.1%) model.
@@
$X^{\prime}=\operatorname{CrossAttn}(E_{t}(X_{q}),\operatorname{SelfAttn}(E_{v}(X_{v}),\mathbf{q})),\vspace{-2mm}$ (8)
where $E_{t}$ is a text tokenizer, $E_{v}$ is the visual tokenizer, $\mathbf{q}$ is the learnable query token. We utilize a decoder $D$ to transform the student-QFormer output, ensuring consistency and recoverability to the teacher’s feature. This supervision bolsters the effectiveness and performance of the model with fewer frames. Then the distillation objective is defined by:
$L_{\text{distill}}=\operatorname{MSE}(D(X^{\prime}_{\text{student}}),X^{\prime}_{\text{teacher}}).\vspace{-2mm}$ (9)
We carefully design the decoder to be a simple Fully Connected (FC) layer with a layer normalization. We show through an ablation study (Sec. 4.3.2) that this combination works the best. Meanwhile, our FP+QFD 4-frames boosts accuracy by 1.8% compared with FP 4-frames model and the key-frame selected overlaps 56.8% with that of the FP 4-frames model. This helps verify our QFD’s ability to both boost performance and enhance the key-frame selection.
4 Experiments
4.1 Implementation Settings